### v1

@@ 1 Introduction
One discipline of IR is Legal Information Retrieval (LIR). In this case, the document corpus includes judicial decisions, legislation, case law, contracts, jurisprudence, and other legal texts, which are often written in specialized language [3]. Its importance lies not only in supporting legal professionals, researchers, and policymakers but also in enabling citizens to exercise their rights and engage with the legal system, which is essential in democratic societies.
@@ 3 Methodology
This section describes the design of JurisTCU, an IR dataset derived from the "selected jurisprudence" database. We cover the document corpus, the queries, and the query relevance judgments.
@@ 3.2 Queries: data collection, data preparation, and description of the queries
The first group of queries (G1) includes the 50 most used queries. These queries were generated by real users and show their actual interactions with the search tool. They are in keyword format, without qualifiers, punctuation, prepositions, or other non-essential elements, in line with the type of search implemented (BM25). Over time, users have likely adapted to this format, recognizing that entering essential and well-chosen keywords is often enough for a successful search. On average, the queries in this group have 3.5 words.
@@ 3.3 Relevance judgment annotation: qrels generation
Considering these trade-offs, we adopted a two-step process: an initial automated evaluation using LLMs, followed by a manual review by a domain expert to correct the judgments produced by the automated process.
The annotation process starts with a two-stage retriever to select potentially relevant documents for each query. The first stage selects a subset of candidate documents, and the second stage reranks them. Although classical lexical retrievers are usually employed in the first stage, the use of semantic retrievers in this stage has also been discussed [16]. Therefore, in the first stage we used both retrievers—a lexical one (BM25) and a semantic one (BERT-based [^6] https://huggingface.co/rufimelo/Legal-BERTimbau-sts-large-ma-v3). The top 300 documents returned by each retriever are combined and then sent to the second stage, where the 600 documents are reranked using a T5-based model [^7] https://huggingface.co/unicamp-dl/mt5-3B-mmarco-en-pt. The top 10 ranked results are selected for annotation, along with 5 randomly chosen documents from the top 1 000 documents retrieved by BM25.
Multi-stage retrievers are common in IR due to their strong performance [16]. Therefore, this process selects 10 documents with a high likelihood to be relevant and 5 with a high likelihood to be irrelevant. These 15 documents are sent to an LLM for automatic annotation, where relevance scores are assigned on a scale from 0 (irrelevant) to 3 (highly relevant). The evaluation used the GPT-4 model with the prompt in Figure 8 in Appendix A. As a result, each query is expected to have 15 annotated documents with varying degrees of relevance.
While some researchers suggest that LLMs could potentially replace human evaluation, the extent of their universal applicability remains a subject of ongoing debate [17]. To address this, a domain expert reviewed the automatically generated relevance judgments. The suggested corrections were implemented, resulting in the final qrels. Figure 3 summarizes the annotation pipeline.
@@ 4.1.1 Lexical retrieval with document expansion
  • docT5query [20]: This method uses a T5-based model to generate synthetic queries that could be used to retrieve a given document. In this experiment, we used a model trained in Portuguese [^8] https://huggingface.co/doc2query/msmarco-portuguese-mt5-base-v1 to generate five synthetic queries from the SUMMARY, which were then appended to the document;
@@ 5 Conclusion
For future work, it is essential for the Portuguese-language IR community to have a diverse collection of datasets available, similar to the English-language IR community. Therefore, we plan to release new annotated datasets to support further research and development in this field.

### v2

@@ 1 Introduction
One subfield of IR is Legal Information Retrieval (LIR). In this case, the document corpus includes judicial decisions, legislation, case law, contracts, jurisprudence, and other legal texts, which are often written in specialized language [3]. Its importance lies not only in supporting legal professionals, researchers, and policymakers but also in enabling citizens to exercise their rights and engage with the legal system, which is essential in democratic societies.
@@ 2 Related work
In addition to existing datasets, it is worth noting that LLMs have been extensively explored and adapted for legal applications. Several models have been trained or fine-tuned specifically for the legal domain, such as SaulLM-7B, SaulLM-54B, and SaulLM-141B [13, 14]; and BERT-like models [15]. To evaluate the performance of LLMs in legal contexts, several benchmarks have been proposed across jurisdictions and languages, such as LawBench [16], LexGLUE [17], and LEXTREME [18]. Some professional legal exams, such as the U.S. Uniform Bar Exam [19], and the Brazilian Bar exam [20, 21] have also been used to assess the legal reasoning capabilities of LLMs. In retrieval-augmented contexts, LexRAG evaluates multi-turn legal consultation using RAG techniques [22], while LegalBench-RAG focuses specifically on precision in legal text retrieval [23].
@@ 3 Methodology
This section describes the design of JurisTCU, an IR dataset derived from the "selected jurisprudence" database. First, we provide a brief background on this database. Then, we cover the document corpus, the queries, and the query relevance judgments.
3.1 A brief overview of the TCU’s role and its jurisprudential databases
The JurisTCU dataset was built using documents produced by TCU, which is responsible for overseeing the use of public resources in Brazil. When potential irregularities or legal issues are identified during audits, the matter is judged by a collegiate body composed of ministers. The decision is reported in an official document called an Acórdão, which is authored by a designated minister known as the rapporteur.
Each Acórdão generally includes a summary, which briefly states the decision; a report, which outlines the facts, evidence, and procedural background; and the vote, which presents the reasoning and legal interpretation adopted by the rapporteur and approved by the plenary.
These documents can be extensive, often spanning dozens of pages, and represent how the TCU interprets and applies legal norms in concrete situations. With over 500 000 documents dating back to 1992 and made publicly available, they serve as precedents - a legal decision or form of proceeding serving as an authoritative rule or pattern in future similar cases.
These precedents collectively form the jurisprudence of the TCU. While not legally binding in the same way as in common law jurisdictions, TCU jurisprudence has normative weight and is frequently cited to support consistent interpretation of administrative law and public finance rules in Brazil.
To facilitate access to its jurisprudence, the TCU maintains a curated collection called "selected jurisprudence", which includes cases deemed to have high precedential value. This subset is the basis for the JurisTCU dataset.
Each document in the "selected jurisprudence" dataset is a synthesis of a prior decision, typically derived from an Acórdão. These documents have a concise statement of the legal understanding extracted form the decision (a summary), and a chunk of text with the original decision that provides textual support for that statement (an excerpt).
@@ 3.3 Queries: data collection, data preparation, and description of the queries
The first group of queries (G1) includes the 50 most used queries. These queries were generated by real users and show their actual interactions with the search tool. They are in keyword format, without qualifiers, punctuation, prepositions, or other non-essential elements, in line with the type of search method used (BM25). Over time, users have likely adapted to this format, recognizing that entering essential and well-chosen keywords is often enough for a successful search. On average, the queries in this group have 3.5 words.
@@ 3.4 Relevance judgment annotation: qrels generation
Considering these trade-offs, we adopted a two-step process: an initial automated evaluation using LLMs, followed by a manual review by a domain expert to correct the judgments produced by the automated process. This approach aligns with the use of multi-stage retrievers, which are commonly employed in IR due to their strong performance [27] and accelerates the creation of qrels while still preserving a degree of human oversight
The annotation process starts with a two-stage retriever to select potentially relevant documents for each query. The first stage selects a subset of candidate documents, and the second stage reranks them. Although classical lexical retrievers are usually employed in the first stage, the use of semantic retrievers in this stage has also been discussed [27]. Therefore, in the first stage we used both retrievers—a lexical one (BM25) and a semantic one (rufimelo/Legal-BERTimbau-sts-large-ma-v3 BERT-based model [^6] https://huggingface.co/rufimelo/Legal-BERTimbau-sts-large-ma-v3). Considering all queries, the overlap between the top 300 documents returned by both retrievers ranged from 1.3% to 44.6%, averaging 15.89% across the 150 queries. The top 300 documents returned by each retriever are combined and then sent to the second stage, where the 600 documents are reranked using a T5-based model (unicamp-dl/mt5-3B-mmarco-en-pt [^7] https://huggingface.co/unicamp-dl/mt5-3B-mmarco-en-pt). The top 10 ranked results are selected for annotation, along with 5 randomly chosen documents from the top 1 000 documents retrieved by BM25.
The purpose of selecting 5 random documents from the BM25 output is to provide documents with different levels of relevance for annotation, as these documents are highly likely to be irrelevant. This approach is commonly used when training ranking models (e.g., [28]), where both relevant and non-relevant examples are needed. Non-relevant documents can be selected randomly from the corpus (random negatives) or from retriever outputs (hard negatives) [29, 30]. Hard negatives, like those from BM25, share some terms with the query but are not truly relevant, making them more useful for evaluating retrieval methods.
These 15 documents are sent to an LLM for automatic annotation, where relevance scores are assigned on a scale from 0 (irrelevant) to 3 (highly relevant). The evaluation used the GPT-4 model with the prompt in Figure 8 in Appendix A. As a result, each query is expected to have 15 annotated documents with varying degrees of relevance.
While some researchers suggest that LLMs could potentially replace human evaluation, the extent of their universal applicability remains a subject of ongoing debate [31]. To address this, a domain expert reviewed the automatically generated relevance judgments to correct them if necessary, resulting in the final qrels. In the case of JurisTCU, the expert agreed with the automatic evaluation. We believe two factors contributed to this agreement. First, LLMs used in legal contexts have often produced results that outperform those of human lawyers [32]. Second, expert validation of pre-generated results can introduce confirmation bias. Unless the output is highly inconsistent (e.g., a clearly irrelevant document marked as highly relevant), experts tend to agree with the judgments they are reviewing.
Figure 3 summarizes the annotation pipeline.
@@ 4.1.1 Lexical retrieval with document expansion
  • docT5query [35]: This method uses a T5-based model to generate synthetic queries that could be used to retrieve a given document. In this experiment, we used a model trained in Portuguese (doc2query/msmarco-portuguese-mt5-base-v1888https://huggingface.co/doc2query/msmarco-portuguese-mt5-base-v1) to generate five synthetic queries from the SUMMARY, which were then appended to the document;
@@ 5 Conclusion
For future work, it is essential for the Portuguese-language IR community to have a diverse collection of datasets available, similar to the English-language IR community. Therefore, we plan to release new annotated datasets to support further research and development in this field. Another possibility is the manual annotation of the JurisTCU dataset, which would enable progress in studies comparing manual and automatic relevance judgments in Portuguese.